OpenStudy (anonymous):

Find two numbers who's difference is 10 and who's product is a minimum? can you walk me through the steps as well?

OpenStudy (mathmale):

This problem involves both algebra and calculus. Choose a letter to represent each of the two numbers. For example, choose x and y to be the two numbers. Given that the 2 numbers are x and y, how would you write a formula for the "difference" of x and y? And what is that "difference" equal to?

OpenStudy (anonymous):

\[x-y=10\]

OpenStudy (anonymous):

?

OpenStudy (anonymous):

that makes \(x=y+10\) so the product is \(y\times (y+10)\) or you could use \(x(x+10)\) as the product you certainly don't need calculus to find the minimum of the quadratic \[x^2+10x\]

OpenStudy (anonymous):

matter of fact you don't need algebra at all

OpenStudy (mathmale):

Yes. x-y = 10. Also, the area is A = x * y. You'll need to solve x-y=10 for x and then substitute the result into A = x*y. Would you please do that? satellite73 has already gotten there. Think of how you could find the x-value at which you have the max. area

OpenStudy (anonymous):

minimum is what you are looking for right?

OpenStudy (anonymous):

you don't really need algebra at all since you cannot tell the numbers apart, i.e. it is symmetric in \(x\) and \(y\) the minimum occurs when they are equidistant from 0, i.e. one is 5 and the other is -5 but you can use algebra to find the vertex of \(A=x^2+10x\) if you like

OpenStudy (anonymous):

\[A=(x^2+10x)*x\]

OpenStudy (anonymous):

?

OpenStudy (mathmale):

I agree with satellite73's \[A=x^2+10x\]

OpenStudy (anonymous):

so just\[x^2+10x\]

OpenStudy (mathmale):

Erob: graph this: A-x^2 + 10x. You could then find the x-value at which you have a minimum just by looking at the graph.

OpenStudy (mathmale):

Use the Draw utility, please.

OpenStudy (anonymous):

can just one person explain because im getting confused on which step and where im suppose to put things? and start over?

OpenStudy (mathmale):

All right. Hold a minute. I'm tieing up another matter.

OpenStudy (anonymous):

k

OpenStudy (mathmale):

Find two numbers who's difference is 10 and who's product is a minimum? can you walk me through the steps as well? choosing x and y to represent the 2 numbers, x-y=10 is the "difference." Thus, x=y+10. Do you believe that?

OpenStudy (anonymous):

yes

OpenStudy (mathmale):

And you want to MINIMIZE the product, P, of the two numbers, correct?

OpenStudy (anonymous):

yeah

OpenStudy (mathmale):

Good. Then your product, P = x*y, becomes (y+10)*y after we substitute x = y+10. Thus, P(y) = (y+10)*y. OK with this?

OpenStudy (anonymous):

ok got that so y^2+10y

OpenStudy (mathmale):

Yes. Label it. P(y) = y^2 + 10y. Just by looking at this, we can see that the graph is that of a parabola that opens up. Correct?

OpenStudy (anonymous):

yeah

OpenStudy (mathmale):

Therefore, the graph has a minimum value. We have to find that minimum value. Know how to do that?

OpenStudy (anonymous):

no....

OpenStudy (mathmale):

We could either graph the function P(y)=y^2+10y and look to see which y value results in a minimum P, or we could use the formula y=-b/(2a) to find the y-coordinate of the vertex. The vertex represents the minimum in this case. What approach would you prefer?

OpenStudy (anonymous):

I already have the graph so we can do it that way.

OpenStudy (mathmale):

Nice work. Please estimate what y value results in P being at its MINIMUM.

OpenStudy (anonymous):

-24?

OpenStudy (mathmale):

Erob: This seems to be a graph of a function different from ours. Ours is P(y)=1y^2+10y. Are you sure that's what you've graphed?

OpenStudy (anonymous):

Yeah hold on let me check

OpenStudy (anonymous):

yeah it is

OpenStudy (anonymous):

we can just use the equation it's no problem

OpenStudy (mathmale):

I'm much more comfortable with your new graph. Thank you. Notice that this graph represents the PRODUCT of x and y. If you'll go down to the vertex (the minimum), you'll see that the y-value of the vertex is -5 (halfway in between -10 and 0). Agreed?

OpenStudy (anonymous):

ok

OpenStudy (mathmale):

So it appears that our solution is y=-5, x= y + 10 = 5.

OpenStudy (anonymous):

ok so x=5 and y=-5

OpenStudy (mathmale):

Look what happens: If x= 5 and y=-5, their product is -25. if x=6 and y=-4, their product is -24 so it really does appear that we get the smallest product when x=5 and y=-5. OK with this?

OpenStudy (anonymous):

yeah thank for the help

OpenStudy (mathmale):

Happy that we've reached a satisfactory solution! Hope to work with you again. Take care, Erob! bye.